OpenStudy (anonymous):

When the consumer price index decreases from 250 to 150, prices decrease by100%

jimthompson5910 (jim_thompson5910):

I'm not 100% sure on how the CPI works, but my guess is that we have a change of 250-150 = 100 this change of 100 is divided by the initial amount 250 to get 100/250 = 10/25 = 2/5 = 0.4 = 40% Going from 250 to 150 means we have had a 40% reduction. So my guess is that the prices (on average) have gone down 40%.